40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12979     Entry time: Wed May 10 01:56:06 2017     In reply to: 12976     Reply to this: 12981   12983
Author: gautam 
Type: Update 
Category: General 
Subject: MICH NB - OL coupling 

Last night, I tried to estimate the contribution of OL feedback signal to the MICH length error signal.

In order to do so, I took a swept sine measurement with a few points between 50 Hz and 500 Hz. The transfer function between C1:LSC-MICH_OUT_DQ and the Oplev Servo Output point (e.g. C1:SUS-BS_OL_PIT_OUT etc) was measured. I played around with the excitation amplitude till I got coherence > 0.9 for the TF measurement, while making sure I wasn't driving the Oplev error point too hard that side-lobes began to show up in the MICH control signal spectrum.

The Oplev control signal is not DQ-ed. So I locked the DRMI again and downloaded the 16k data "live" for ~5min stretch using cdsutils.getdata on the workstation. The Oplev error point is DQ-ed at 2k, but I found that the excitation amplitude needed for good SNR at the error point drove the servo to the limiter value of 2000cts - so I decided to use the control signal instead. Knowing the transfer function from the Oplev *_OUT* channel to C1:LSC-MICH_IN1_DQ, I backed out the coupling - the transfer function was only measured between 50 Hz and 500 Hz, and no extrapolation is done, so the estimation is only really valid in this range, which looks like where it is important anyways (see Attachment #2, contributions from ITMX, ITMY and BS PIT and YAW servos added in quadrature).

I was also looking at the Oplev servo shapes and noticed that they are different for the ITMs and the BS (Attachment #1). Specifically, for the ITM Oplevs, an "ELP15" is used to do the roll-off while an "ELP35" is employed in the BS servo (though an ELP35 also exists in the ITM Oplev filter banks). I got lost in an elog search for when these were tuned, but I guess the principles outlined in this elog still hold and can serve as a guideline for Oplev loop tweaking.

Coil driver noise estimation to follow

Quote:

I think the most important next two items to budget are the optical lever noise, and the coil driver noise. The coil driver noise is dominated at the moment by the DAC noise since we're operating with the dewhitening filters turned off.

GV 10 May 12:30pm: I've uploaded another copy of the NB (Attachment #3) with the contributions from the ITMs and BS separated. Looks like below 100Hz, the BS coupling dominates, while the hump/plateau around 350Hz is coming from ITMX.

Attachment 1: OL_BS_ITM_comp.pdf  16 kB  Uploaded Wed May 10 03:09:18 2017  | Hide | Hide all
OL_BS_ITM_comp.pdf
Attachment 2: C1NB_disp_40m_MICH_NB_8_May_2017.pdf  785 kB  Uploaded Wed May 10 03:15:32 2017  | Hide | Hide all
C1NB_disp_40m_MICH_NB_8_May_2017.pdf
Attachment 3: C1NB_disp_40m_MICH_NB_10_May_2017.pdf  817 kB  Uploaded Wed May 10 13:29:44 2017  | Hide | Hide all
C1NB_disp_40m_MICH_NB_10_May_2017.pdf
ELOG V3.1.3-